When 1 g of water at 373K is converted into steam at the same temperature, the appriximate change in volume is?

Grade:10

1 Answers

Chetan Mandayam Nayakar
13 years ago

volume of steam=nRT/P, n=number of moles=1/18, T=373, P=pressure, volume of water=1ml, subtracting the two you get the change